OpenStudy (anonymous):

6 x + 9 / 15 x^2 all over 16 x - 12 / 10 x^4

OpenStudy (anonymous):

2x+3/2(4x-3) correct?

OpenStudy (anonymous):

You are missing a component on your numerator

OpenStudy (anonymous):

x squared in front of the 2x + 3 right? x^2(2x+3)?

OpenStudy (anonymous):

if you look at your original equation, you have x^4 components and x^2 so you can't have a solution that has the same coefficients on top and bottom

OpenStudy (anonymous):

if you didn't have an x^4 in your equation, and instead had an x^2; then your solution would be correct :)

OpenStudy (anonymous):

so yes?

OpenStudy (anonymous):

You have \[\left( \frac{ 6x+9 }{ 15x ^{2} } \right) \div \left( \frac{ 16x-12 }{ 10x ^{4} } \right)\] which can be rewritten as \[\frac{ (6x+9)*(10x ^{4}) }{ (16x-12)*(15x ^{2}) }\] if you expand this out and factorise you get; \[\frac{ 30x ^{4}(2x+3) }{ 30x ^{2}(8x-6) }\] Cancel out, you get; \[\frac{ x ^{2}(2x+3) }{ (8x-6) }\] Do you understand?